a normal force of 35 N acts over an area of 12.5cm2 what is the pressure exerted by the force?


1
Expert's answer
2022-06-01T12:49:09-0400
"p=\frac{F}{A}=\frac{35\:\rm N}{12.5*10^{-4}\:\rm m^2}=28000\:\rm Pa"